# Large Language Models for Anomaly and Out-of-Distribution Detection: A Survey

### v2

@@ 1 Introduction
Large Language Models (LLMs), such as GPT-4 Achiam et al. (2023) and LLaMA Touvron et al. (2023), have demonstrated remarkable capabilities in language comprehension and generation. To further harness the potential of LLMs beyond text data, there is also a growing interest in extending them to multi-modal tasks such as vision-language understanding and generation Wang et al. (2024), evolving them into Multimodal LLMs (MLLMs) Yin et al. (2023). Given the zero- and few-shot reasoning capabilities of LLMs and MLLMs, researchers try to apply these models to anomaly and out-of-distribution (OOD) detection, as illustrated in Figure 1, yielding promising detection results.
@@
Remarkably, the emergence of LLMs has fundamentally changed the learning paradigm in this field, compared to previous statistical and machine learning methods. In the meantime, while leveraging LLMs to solve the problem of anomaly and OOD detection has drawn much attention, this field remains underexplored, highlighting the need for a comprehensive survey to analyze the emerging challenges and systematically review the rapidly expanding works. Recently, Salehi et al. (2021) and Yang et al. (2024a) present unified frameworks for OOD detection but do not delve into the utilization of LLMs. While Su et al. (2024) review some small-sized language models for forecasting and anomaly detection, they neither cover the usage of LLMs with emergent abilities nor discuss OOD detection. A recent survey by Miyai et al. (2024a) summarizes works on anomaly and OOD detection in vision using vision-language models but neglects other data modalities. Therefore, we aim to conduct a systematic survey that covers both anomaly and OOD detection tasks across various data domains, concentrating on how LLMs are used in existing works.
@@
In this survey, we propose a novel taxonomy that focuses on how LLMs can profoundly impact anomaly and OOD detection in two fundamental ways, as illustrated in Figure 4: ❶ LLMs for Detection (§3): We provide a detailed review of existing methods that leverage LLMs as detectors for identifying anomalies and OOD instances, categorizing these methods into distinct pipelines; and ❷LLMs for Generation (§4): We also review methods that utilize LLMs’ emergent abilities, advanced semantic understanding, and vast knowledge to generate augmented data and explanations, thereby enhancing detection models with more robust and interpretable outputs. At the end (§5), we also outline the challenges and future research directions, in order to provide a better understanding of anomaly and OOD detection in the era of LLMs and shed light on the following research.
@@ 3.2 Contrasting-based Detection
We further categorize contrasting-based detection methods into two main classes depending on whether there exists additional training and fine-tuning.
@@ 3.2.1 Detection without LLM Tuning
    For OOD detection, to address the challenges of using only in-distribution (ID) class information while avoiding the matching of OOD inputs to irrelevant ID classes, one notable approach is the Maximum Concept Matching (MCM) framework proposed by Ming et al. (2022). This method is not limited to CLIP and can be generally applicable to other pre-trained models that promote multi-modal feature alignment. They view the textual embeddings of ID classes as a collection of concept prototypes and define the maximum concept matching (MCM) score based on the cosine similarity between the image feature and the textual feature. Following the idea of MCM, several subsequent works focus on improving OOD detection results by either adding a local MCM score or modifying weights in the original MCM framework, such as Miyai et al. (2023) and Li et al. (2024c).
@@ 4 LLMs for Generation
In this section, we review methods that leverage LLMs as generative tools for enhancing anomaly and OOD detection. LLMs use their extensive pre-trained knowledge to generate augmented data, such as embeddings, pseudo labels, and textual descriptions, which improve detection performance. Furthermore, due to their ability to understand and generate human-like text, LLMs have been explored for providing insightful explanations and analyses of detection results, aiding in interpretation, planning, and decision-making. These methods are classified into two main approaches: Augmentation-centric Generation and Explanation-centric Generation.
@@ 4.1.2 Pseudo Label-based Augmentation
The emergent capabilities of LLMs provide a promising approach for generating high-quality synthetic datasets, including pseudo labels for OOD samples. A significant challenge in OOD detection is the lack of labeled OOD data, which can limit model performance. Traditionally, obtaining OOD labels requires extensive human effort, but LLMs can mitigate this by generating pseudo-OOD labels through carefully designed prompts. These pseudo labels augment existing ID data and enhance the distinction between ID and OOD samples.
@@ 4.2 Explanation-centric Generation
For example, Holmes-VAD Zhang et al. (2024b) trains a lightweight temporal sampler to select frames with high anomaly scores and uses an LLM to generate detailed explanations, providing clear insights into the detected anomalies. VAD-LLaMA Lv and Sun (2024) generates instruction-tuning data to train the projection layer of Video-LLaMA, enabling more comprehensive explanations of anomalies. AnomalyRuler Yang et al. (2024b) employs a rule-based reasoning strategy with few-normal-shot prompting, providing interpretable, rule-driven explanations that can quickly adapt to various video anomaly detection scenarios.
@@ 6 Conclusion
In this survey, we examined the use of Large Language Models (LLMs) and multimodal LLMs (MLLMs) in anomaly and out-of-distribution (OOD) detection. We introduced a novel taxonomy categorizing methods into two approaches based on the role of LLMs in the architectures: detection, and generation. This taxonomy clarifies how LLMs can augment data, detect anomalies or OOD, and build explainable systems. We also discussed limitations and future research directions, aiming to highlight advancements and challenges in the field of anomaly or OOD detection and encourage further progress.
@@
Appendix A A Collection of Papers on LLMs for Anomaly and OOD Detection
We provide a taxonomy of collected papers on LLMs for anomaly and OOD detection, which is summarized in 4.
{forest}
CLIPScope  Fu et al. (2024), WinCLIP  Jeong et al. (2023),
AnoCLIP Deng et al. (2023), CLIP-AD Chen et al. (2023b), MCM  Ming et al. (2022),
 Miyai et al. (2023), SETAR Li et al. (2024c), leaf, text width=38em ] ] [ w/ Tuning,text width=8.3em,middle [ LoCoOp Miyai et al. (2024c) AnomalyCLIP Zhou et al. (2024), InCTRL
 Zhu and Pang (2024), MVFA  Huang et al. (2024a), ID-like Bai et al. (2024),
NegPrompt Li et al. (2024b), CLIPN Wang et al. (2023),
 Liu et al. (2024a),  Zhang et al. (2024a), leaf, text width=38.2em ] ] [ Psuedo Label,text width = 8.2em,middle [ EOE Cao et al. (2024), PCC Huang et al. (2024b),
TOE Park et al. (2023), CoNAL Xu et al. (2023), leaf, text width=38.2em ] ] [ Textual Description,text width = 8.2em,middle [ TagFog  Chen et al. (2024), ALFA Zhu et al. (2024),  Dai et al. (2023), leaf, text width=38.2em ] ] ] [ Explanation,middle,text width=5.8em [ Holmes-VAD Zhang et al. (2024b), AnomalyRuler Yang et al. (2024b),
Appendix B LLM-based Anomaly and OOD Detection: Strengths and Challenges
B.1 Strengths
Zero-shot/Few-shot:
@@ B.1 Strengths
Traditional anomaly and OOD detection methods require extensive training on well-defined normal and in-distribution datasets, which can be both time-consuming and computationally expensive. In contrast, LLMs can perform zero-shot or few-shot reasoning or learning, producing detection results without needing large-scale training Kojima et al. (2022). This allows researchers to bypass the complex process of data collection and model training, enabling faster anomaly and OOD detection.
Explainability and Interpretability:
@@
LLMs possess strong reasoning abilities that can contribute to building explainable systems for anomaly and OOD detection. Traditional methods often rely on scores that offer little insight into the detection process. In contrast, LLMs can provide detailed, interpretable explanations for their detection results, offering valuable insights for future actions. Furthermore, LLMs can be integrated as agents within a system, assisting in planning the next steps when an anomaly or OOD event is detected.
B.2 Challenges
Computational Efficiency and Token Limits:
@@ B.2 Challenges
A major concern when leveraging LLMs for anomaly and OOD detection is computational inefficiency. Applying LLMs in these tasks often requires complex reasoning which can lead to significant computational overhead. Additionally, many LLMs have input token limits, making it impossible to feed large amounts of data directly into the model. To address this, researchers must carefully design architectures that allow LLMs to process the data effectively. For instance, techniques like Retrieval-Augmented Generation have been used to retrieve the most relevant data to avoid token limit issues Liu et al. (2024c). Moreover, methods such as model pruning and knowledge distillation should be considered to reduce computational costs while maintaining high accuracy.
Domain Knowledge:
@@
While LLMs are trained on vast and diverse datasets, they may lack specific domain expertise needed for certain anomaly and OOD detection tasks. To enhance performance in these specialized domains, incorporating domain knowledge into the LLMs is crucial. One strategy is injecting domain knowledge into prompts to guide the LLM’s understanding. Another approach involves using adapters or fine-tuning to better tailor the LLMs to domain-specific problems, ensuring they perform well in specialized tasks.
Hallucination and Trustworthiness:
@@
LLMs can sometimes produce inaccurate or fabricated information, a phenomenon known as hallucination. In the context of anomaly and OOD detection, hallucinations pose a significant risk, potentially leading to incorrect or misleading results. To mitigate this, researchers need to work on reducing hallucination rates and improving the trustworthiness of the model. Manual checks may still be necessary in critical applications, as LLMs should be seen as assistants rather than sole decision-makers.
Appendix C Anomaly Detection Research Roadmap
@@
C.1 Traditional Methods
Density-based Methods.
Density-based methods model the distribution of normal data and detect anomalies by evaluating how well a sample fits this modeled distribution. The underlying assumption is that normal data is more likely to have a higher likelihood under the distribution, while anomalous data will have a lower likelihood.
@@ C.1 Traditional Methods
Reconstruction-based Methods.
Reconstruc-
@@
Distance-based Methods.
Distance-based methods detect anomalies by measuring the distance between test samples and reference points, such as class prototypes or centroids. Anomalies are expected to be further from these reference points than normal samples Tian et al. (2014).
Classification-based Methods.
Classification-based methods treat anomaly detection as a supervised learning problem. In one-class classification (OCC) Tax (2002), the goal is to learn a decision boundary that encompasses the normal data, with any data points outside this boundary being classified as anomalies.
@@
C.2 LLM-based Anomaly Detection
@@
C.3 Comparison: Traditional vs. LLM-based Anomaly Detection
@@
  • Assumptions: Traditional methods often rely on predefined assumptions about the data distribution, such as the parametric forms used in density-based methods. In contrast, LLM-based approaches require fewer assumptions and are better equipped to generalize across a variety of tasks.
@@
  • Explainability: Traditional methods tend to offer high interpretability but lack the ability to explain their decisions in natural language. LLM-based approaches can not only detect anomalies but also provide natural language explanations, which improves transparency and trustworthiness.
@@
  • Generalization: LLM-based methods are highly adaptable, capable of processing different types of data and tasks, such as text and images. In contrast, traditional methods often need custom architectures tailored to the specific data modality.
@@
Appendix D Out-of-Distribution Detection Research Roadmap

### v3

@@ 1 Introduction
Large Language Models, such as GPT-4 Achiam et al. (2023) and LLaMA Touvron et al. (2023), have demonstrated remarkable capabilities in language comprehension and generation. To further harness the potential of LLMs beyond text data, there is also a growing interest in extending them to multi-modal tasks such as vision-language understanding and generation Wang et al. (2024), evolving them into Multimodal LLMs (MLLMs) Yin et al. (2023). Given the zero- and few-shot reasoning capabilities of LLMs and MLLMs, researchers try to apply these models to anomaly and out-of-distribution (OOD) detection, as illustrated in Figure 1, yielding promising detection results.
@@
Remarkably, the emergence of LLMs has fundamentally changed the learning paradigm in this field. In the meantime, while leveraging LLMs to solve the problem of anomaly and OOD detection has drawn much attention, this field remains underexplored, highlighting the need for a comprehensive survey to analyze the emerging challenges and systematically review the rapidly expanding works. Recently, Salehi et al. (2021) and Yang et al. (2024a) present unified frameworks for OOD detection but do not delve into the utilization of LLMs. While Su et al. (2024) review some small-sized language models for forecasting and anomaly detection, they neither cover the usage of LLMs with emergent abilities nor discuss OOD detection. A recent survey by Miyai et al. (2024a) summarizes works on anomaly and OOD detection in vision using vision-language models but neglects other data modalities. Therefore, we aim to conduct a systematic survey that covers both anomaly and OOD detection across various data domains, concentrating on how LLMs are used in existing works.
@@
In this survey, we propose a novel taxonomy that focuses on how LLMs can profoundly impact anomaly and OOD detection in two fundamental ways, as illustrated in Figure 2: ❶ LLMs for Detection (§3): We provide a detailed review of existing methods that leverage LLMs as detectors for identifying anomalies and OOD instances; and ❷LLMs for Generation (§4): We also review methods that utilize LLMs’ emergent abilities, advanced semantic understanding, and vast knowledge to generate augmented data and explanations. At the end (§5 and §6), we also summarize widely used datasets and outline future research directions, in order to provide a better understanding of anomaly and OOD detection in the era of LLMs and shed light on the following research.
{forest}
CLIPScope  Fu et al. (2024), WinCLIP  Jeong et al. (2023),
AnoCLIP Deng et al. (2023), CLIP-AD Chen et al. (2023b), MCM  Ming et al. (2022),
 Miyai et al. (2023), SETAR Li et al. (2024d), leaf, text width=38em ] ] [ w/ Tuning,text width=8.3em,middle [ LoCoOp Miyai et al. (2024c) AnomalyCLIP Zhou et al. (2024), InCTRL
 Zhu and Pang (2024), MVFA  Huang et al. (2024a), ID-like Bai et al. (2024),
NegPrompt Li et al. (2024b), CLIPN Wang et al. (2023),
 Liu et al. (2024a),  Zhang et al. (2024a), leaf, text width=38.2em ] ] [ Pseudo Label,text width = 8.2em,middle [ EOE Cao et al. (2024), PCC Huang et al. (2024b),
TOE Park et al. (2023), CoNAL Xu et al. (2023), leaf, text width=38.2em ] ] [ Textual Description,text width = 8.2em,middle [ TagFog  Chen et al. (2024), ALFA Zhu et al. (2024),  Dai et al. (2023), leaf, text width=38.2em ] ] ] [ Explanation,middle,text width=5.8em [ Holmes-VAD Zhang et al. (2024b), AnomalyRuler Yang et al. (2024c),
@@ 3.2 Contrasting-based Detection
We further categorize contrasting-based detection methods into two classes depending on whether there exists additional training and fine-tuning.
@@ 3.2.1 Detection without LLM Tuning
    For OOD detection, to address the challenges of using only in-distribution (ID) class information while avoiding the matching of OOD inputs to irrelevant ID classes, one notable approach is the Maximum Concept Matching (MCM) framework proposed by Ming et al. (2022). This method is not limited to CLIP and can be generally applicable to other pre-trained models that promote multi-modal feature alignment. They view the textual embeddings of ID classes as a collection of concept prototypes and define the maximum concept matching (MCM) score based on the cosine similarity between the image feature and the textual feature. Following the idea of MCM, several subsequent works focus on improving OOD detection results by either adding a local MCM score or modifying weights in the original MCM framework, such as Miyai et al. (2023) and Li et al. (2024d).
@@ 4 LLMs for Generation
In this section, we review methods that leverage LLMs as generative tools for enhancing anomaly and OOD detection. LLMs use their extensive pre-trained knowledge to generate augmented data, such as embeddings, pseudo labels, and textual descriptions, improving detection performance Li et al. (2024c); Ding et al. (2024). Furthermore, due to their ability to understand and generate human-like text, LLMs have been explored for providing insightful explanations and analyses of detection results, aiding in interpretation, planning, and decision-making. These methods are classified into two main approaches: ❶ Augmentation-centric Generation and ❷ Explanation-centric Generation.
@@ 4.1.2 Pseudo Label-based Augmentation
The emergent capabilities of LLMs provide a promising approach for generating high-quality synthetic datasets, including pseudo labels for OOD samples. A significant challenge in OOD detection is the lack of labeled OOD data, which can limit model performance. Traditionally, obtaining OOD labels requires extensive human effort, but LLMs can mitigate this by generating pseudo-OOD labels through carefully designed prompts.
@@ 4.2 Explanation-centric Generation
For example, Holmes-VAD Zhang et al. (2024b) trains a lightweight temporal sampler to select frames with high anomaly scores and uses an LLM to generate detailed explanations, providing clear insights into the detected anomalies. VAD-LLaMA Lv and Sun (2024) generates instruction-tuning data to train the projection layer of Video-LLaMA, enabling more comprehensive explanations of anomalies. AnomalyRuler Yang et al. (2024c) employs a rule-based reasoning strategy with few-normal-shot prompting, providing interpretable, rule-driven explanations that can quickly adapt to various video anomaly detection scenarios.
@@
5 Evaluation Datasets
In this section, we introduce commonly used datasets across multiple modalities in Table 1, including images, videos, text, and time series, that serve as benchmarks for anomaly and OOD detection We also introduce widely used evaluation metrics in Appendix E.
Task Type Datasets
Images Anomaly MVTec-AD Bergmann et al. (2019)
Anomaly VisA Zou et al. (2022)
Anomaly Chest X-ray Kermany et al. (2018)
Anomaly Head CT Felipe (2018)
Anomaly LOCO Bergmann et al. (2022)
OOD ImageNet Deng et al. (2009)
OOD iNaturalist Van Horn et al. (2018)
OOD Places Zhou et al. (2017)
OOD SUN Xiao et al. (2010)
OOD Texture Cimpoi et al. (2014)
Videos Anomaly UCF-Crime Sultani et al. (2018)
Anomaly XD-Violence Wu et al. (2020)
Anomaly VAD-Intruct50k Zhang et al. (2024b)
Texts Anomaly NLP-AD Li et al. (2024e)
Anomaly AD-NLP Bejan et al. (2023)
OOD 20NG Lang (1995)
OOD SST-2 Socher et al. (2013)
OOD CLINC150 Repository (2020)
Times Series Anomaly ODDS Rayana (2016)
Anomaly Yahoo Paparrizos et al. (2022)
Anomaly ECG Paparrizos et al. (2022)
Anomaly SVDB Paparrizos et al. (2022)
Anomaly IOPS Paparrizos et al. (2022)
Table 1: A summary of datasets for Anomaly and Out-of-Distribution (OOD) Detection by Modality.
Images. Industrial anomaly detection is commonly evaluated using MVTec-AD Bergmann et al. (2019) and VisA Zou et al. (2022), which provide benchmarks for defect detection and localization. In the medical domain, datasets such as Chest X-ray Kermany et al. (2018) and Head CT Felipe (2018) are used for detecting abnormalities in clinical imaging. Logical anomaly detection is assessed with MVTec LOCO Bergmann et al. (2022). For OOD detection, which focuses on semantic shifts, large-scale datasets such as ImageNet Deng et al. (2009), iNaturalist Van Horn et al. (2018), Places Zhou et al. (2017), and SUN Xiao et al. (2010) are widely used.
Videos. UCF-Crime Sultani et al. (2018) and XD-Violence Wu et al. (2020) are widely used for detecting anomalous activities in surveillance footage. The VAD-Instruct50k dataset Zhang et al. (2024b) introduces a large-scale multimodal benchmark for video anomaly detection.
Texts. Textual anomaly and OOD detection have gained attention with the rise of LLMs. NLP-AD Li et al. (2024e) and AD-NLP Bejan et al. (2023) provide benchmarks for detecting anomalous patterns in textual data. Recent work, AD-LLM Yang et al. (2024b), evaluates LLM performance on text anomaly detection, demonstrating their superior effectiveness in this task. Textual OOD detection is commonly assessed using CLINC150 Repository (2020), which partitions intent categories into ID and OOD classes. Additional benchmarks, such as 20 Newsgroups (20NG) Lang (1995) and SST-2 Socher et al. (2013), are used for OOD detection in topic classification and sentiment analysis.
Time Series. The ODDS dataset Rayana (2016) serves as a standard benchmark for time series anomaly deteciton, while Yahoo, ECG, SVDB, and IOPS datasets Paparrizos et al. (2022) focus on detecting anomalies in financial transactions, physiological signals, and system performance.
@@ 7 Conclusion
In this survey, we examine the use of Large Language Models (LLMs) and multimodal LLMs (MLLMs) in anomaly and out-of-distribution (OOD) detection. We introduce a novel taxonomy categorizing methods into two approaches based on the role of LLMs in the architectures: detection, and generation. This taxonomy clarifies how LLMs can augment data, detect anomalies or OOD, and build explainable systems. We also summarize commonly used datasets and discuss future research directions, aiming to highlight advancements and challenges in the field of anomaly or OOD detection and encourage further progress.
@@
Appendix A LLM-based Anomaly and OOD Detection: Strengths and Challenges
A.1 Strengths
@@
Zero-shot/Few-shot: Traditional anomaly and OOD detection methods require extensive training on well-defined normal and in-distribution datasets, which can be both time-consuming and computationally expensive. In contrast, LLMs can perform zero-shot or few-shot reasoning or learning, producing detection results without needing large-scale training Kojima et al. (2022). This allows researchers to bypass the complex process of data collection and model training, enabling faster anomaly and OOD detection.
@@
Explainability and Interpretability: LLMs possess strong reasoning abilities that can contribute to building explainable systems for anomaly and OOD detection. Traditional methods often rely on scores that offer little insight into the detection process. In contrast, LLMs can provide detailed, interpretable explanations for their detection results, offering valuable insights for future actions. Furthermore, LLMs can be integrated as agents within a system, assisting in planning the next steps when an anomaly or OOD event is detected.
A.2 Challenges
@@
Computational Efficiency and Token Limits: A major concern when leveraging LLMs for anomaly and OOD detection is computational inefficiency. Applying LLMs in these tasks often requires complex reasoning which can lead to significant computational overhead. Additionally, many LLMs have input token limits, making it impossible to feed large amounts of data directly into the model. To address this, researchers must carefully design architectures that allow LLMs to process the data effectively. For instance, techniques like Retrieval-Augmented Generation have been used to retrieve the most relevant data to avoid token limit issues Liu et al. (2024c). Moreover, methods such as model pruning and knowledge distillation should be considered to reduce computational costs while maintaining high accuracy.
@@
Domain Knowledge: While LLMs are trained on vast and diverse datasets, they may lack specific domain expertise needed for certain anomaly and OOD detection tasks. To enhance performance in these specialized domains, incorporating domain knowledge into the LLMs is crucial. One strategy is injecting domain knowledge into prompts to guide the LLM’s understanding. Another approach involves using adapters or fine-tuning to better tailor the LLMs to domain-specific problems, ensuring they perform well in specialized tasks.
@@
Hallucination and Trustworthiness: LLMs can sometimes produce inaccurate or fabricated information, a phenomenon known as hallucination. In the context of anomaly and OOD detection, hallucinations pose a significant risk, potentially leading to incorrect or misleading results. To mitigate this, researchers need to work on reducing hallucination rates and improving the trustworthiness of the model. Manual checks may still be necessary in critical applications, as LLMs should be seen as assistants rather than sole decision-makers.
Appendix B Anomaly Detection Research Roadmap
@@
B.1 Traditional Methods
Density-based Methods. Density-based methods model the distribution of normal data and detect anomalies by evaluating how well a sample fits this modeled distribution. The underlying assumption is that normal data is more likely to have a higher likelihood under the distribution, while anomalous data will have a lower likelihood.
@@ B.1 Traditional Methods
Reconstruction-based Methods. Reconstruc-
@@
Distance-based Methods. Distance-based methods detect anomalies by measuring the distance between test samples and reference points, such as class prototypes or centroids. Anomalies are expected to be further from these reference points than normal samples Tian et al. (2014).
Classification-based Methods. Classification-based methods treat anomaly detection as a supervised learning problem. In one-class classification (OCC) Tax (2002), the goal is to learn a decision boundary that encompasses the normal data, with any data points outside this boundary being classified as anomalies.
@@
B.2 LLM-based Anomaly Detection
@@
B.3 Comparison: Traditional vs. LLM-based Anomaly Detection
@@
  • Assumptions: Traditional methods often rely on predefined assumptions about the data distribution, such as the parametric forms used in density-based methods. In contrast, LLM-based approaches are better equipped to generalize across a variety of tasks.
@@
  • Explainability: Traditional methods lack the ability to explain their decisions in natural language. LLM-based approaches can not only detect anomalies but also provide natural language explanations, which improves transparency and trustworthiness for further steps.
@@
  • Generalization: LLM-based methods are highly adaptable, capable of processing different types of data, such as text and images. In contrast, traditional methods often need custom architectures tailored to the specific data modality.
@@
Appendix C Out-of-Distribution Detection Research Roadmap